# The Logical Structure of Physical Laws: A Fixed Point Reconstruction

### v1

@@ 6 Conclusion
To connect the abstract schema to physics, we proposed reading admissibility as completion under constitutive constraints. The key structural lesson is that closure is the invariant content. A package counts as an admissible theory precisely when it is stable under the completion map. This aligns naturally with an invariance-style conception of objectivity. What is ”the theory” is what survives (is closed under) the admissibility transformations one treats as constitutive. In the examples, we instantiated $\mathcal{F}$ via an invariance–symmetry correspondence, and showed how QED and GR arise as least fixed points once one fixes a candidate domain and encodes symmetry, locality in the gauge sense, and a truncation criterion.

### v2

@@ 6 Conclusion
To connect the abstract schema to physics, we proposed reading admissibility as completion under constitutive constraints. The key structural lesson is that closure is the invariant content. A package counts as an admissible theory precisely when it is stable under the completion map. This aligns naturally with an invariance-style conception of objectivity. What is ”the theory” is what survives (is closed under) the admissibility transformations one treats as constitutive. In the examples, we instantiated $\mathcal{F}$ via an invariance–symmetry correspondence, and showed how QED and GR is represented as least fixed points once one fixes a candidate domain and encodes symmetry, locality in the gauge sense, and a truncation criterion.